OpenStudy (anonymous):

a regular octagon is inscribed in a circle with center at 0. Given that the area of the circle is 16pi, find the perimeter of the octagon.

OpenStudy (anonymous):

@Hero please help me

hero (hero):

If the area of the circle is \(16\pi\), then \[A = \pi r^2 = 16 \pi\] so \[\pi r^2 = 16\pi\]\[r^2 = 16\]\[r = 4\]

OpenStudy (anonymous):

thank you

hero (hero):

You still have to find the perimeter of the octagon.

hero (hero):

r is only the radius of the circle.

OpenStudy (anonymous):

I know I'm working on in

OpenStudy (anonymous):

ok idk what I'm doing what is perimeter of an octagon?